# DenseVLM: A Retrieval and Decoupled Alignment Framework for Open-Vocabulary Dense Prediction

### v3

@@ 4.2 Comparison with other VLMs
Quantitative evaluation. We perform a comprehensive quantitative evaluation of dense representations across multiple VLMs on the COCO Panoptic [40] and ADE20K Panoptic [72] datasets. As shown in Tab. 1, while previous methods [4, 11, 69, 55] achieve strong zero-shot image classification performance, their performance in region recognition is notably suboptimal. For instance, EVA-CLIP [55] achieves only a Top-1 box classification accuracy of 44.3% on COCO and 33.0% on ADE20K. Although RegionCLIP [71], which is trained on region-text pairs, shows improved performance on COCO, but fails to generalize to datasets such as ADE20K. Moreover, FineCLIP [25] and CLIPSelf [59], which incorporate self-distillation, achieve higher Mask-T classification accuracy but perform poorly in Mask-S classification. In contrast, our method, DenseVLM, notably outperforms these models, achieving a 4.3% improvement in Top-1 Mask-T accuracy and a 3.9% improvement in Top-1 Mask-S accuracy on COCO, highlighting its effectiveness in dense prediction tasks.
@@ 4.3 Ablation study
In our ablation study, DenseVLM is trained on the unlabeled images from the COCO train2017 split using $\timesA4$0 GPUs and evaluated on the val2017 split. The experiments in Sec. 4.3 utilize the ViT-B/16 model from EVA-CLIP [55] for its superior efficiency and capacity.
Framework design exploration. As shown in Tab. 2, we conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [59] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction while introducing foreground bias issues. In contrast, DenseVLM leverages a more efficient feature-cropping strategy and a decoupled alignment framework to achieve superior performance while substantially reducing both training time and GPU memory usage.
@@
Figure 6: Zero-shot comparisons of models pre-trained on datasets with three different scales. We select three training sets from the SA-1B dataset [28]: 100K, 1.1M, and 5.5M seen samples and perform the zero-shot evaluation on the COCO and ADE20K benchmarks.
@@
To address the limitations of category set annotations, we generate category sets using the generative model NXTP [67] based on images and use DeepSeek-R1 [16] to classify foreground and background categories. As shown in Tab. 4, our model still achieves strong performance.
@@ 4.4 Zero-shot comparisons at different data scales
Table 6: Results and comparisons of different backbones. * indicates the model initialized by OpenCLIP [4].
@@
PACL [50] 31.4 - -
OVSeg [38] 24.8 7.1 11.0
MAFT [23] 29.1 10.1 12.6
SED [62] 31.6 11.4 18.6
SCAN [42] 30.8 10.8 13.2
CAT-Seg+CLIPSelf [59] 29.7 10.1 -
CAT-Seg+FineCLIP [25] 32.4 12.2 -
SAN [64] 27.4 10.0 13.0
@@ 4.5 Extending to different backbones
To demonstrate the generality of our method, we extend it to several backbones, including ViT-B/16, R50x4, and ConvNeXt-B from OpenCLIP [4], as well as ViT-B/16 from EVA-CLIP [55]. From Tab. 6, we observe that CLIPSelf [59] is more effective with ViT-based architectures but is less effective with CNN-based architectures, particularly showing degraded performance with R50x4 compared to the baseline. In contrast, our approach yields consistent performance gains across all architectures, thereby broadening its applicability and potential. Furthermore, DenseVLM narrows the performance gap between powerful models like ViT-L/14 fine-tuned by CLIPSelf and lightweight models, with some lightweight models even outperforming the more powerful ones on the Masks-S metric.
@@ 4.6 Application to open-vocabulary dense tasks
VLDet [39] RN50 32.0 50.6 45.8
F-VLM [29] RN50 28.0 - 39.6
BARON-Cap [58] RN50 33.1 54.8 49.1
CORA [60] RN50 35.1 35.5 35.4
RO-ViT [27] ViT-B/16 30.2 - 41.5
RO-ViT [27] ViT-L/16 33.0 - 47.7
F-ViT+CLIPSelf [59] ViT-B/16 25.4 40.9 36.8
F-ViT+FineCLIP [25] ViT-B/16 29.8 45.9 41.7
F-ViT [59] ViT-B/16 17.5 41.0 34.9
@@
VLDet [39] RN50 21.7 29.8 30.1
BARON-Cap [58] RN50 22.6 27.6 27.6
F-VLM [29] RN50 18.6 - 24.2
COR [60] RN50x4 22.2 - -
RO-ViT [27] ViT-B/16 28.0 - 30.2
F-ViT+CLIPSelf [59] ViT-B/16 10.6 7.6 9.3
F-ViT+FineCLIP [25] ViT-B/16 10.4 8.0 9.5
F-ViT [59] ViT-B/16 11.5 12.3 15.4
@@
Open-vocabulary semantic segmentation. We apply DenseVLM models initialized with OpenAI CLIP [52] to open-vocabulary semantic segmentation, including SAN [64] with a frozen backbone and Cat-Seg [5] with a fine-tuned backbone. The models are trained on COCO-Stuff [1] and evaluated on the ADE20K [73] (ADE-150 and ADE-847) and PASCAL Context [49] (PC-459) datasets, using the mean Intersection-over-Union (mIoU) metric. As shown in Tab. 7, DenseVLM consistently improves performance across all evaluation datasets, further enhancing the state-of-the-art performance.
Open-vocabulary object detection. Building on previous work [59], we construct open-vocabulary object detectors using the F-ViT architecture, a two-stage detector built on frozen ViTs from EVA-CLIP [55]. As shown in Tab. 8, we evaluate performance on the OV-COCO [3] benchmark by reporting box AP at IoU 0.5 for base, novel, and overall categories ($_{50}^{\mathrm{novel}}, AP$$_{50}^{\mathrm{base}}$, AP50), and on OV-LVIS [17] benchmark by mean AP for rare (mAPr), common (mAPc), and all categories (mAP). Replacing the frozen EVA-CLIP ViTs with DenseVLM models leads to clear performance improvements on both benchmarks, achieving competitive performance relative to previous methods.
@@ Acknowledgments
  • In Sec. A, we detail the limitations and broader impact of DenseVLM.
  • In Sec. B, we present the training details.
  • In Sec. C, we offer supplementary experiments on input image sizes, region proposals, backbones, category sets and threshold of region denoising for training VLMs.
  • In Sec. D, we show visualizations, including confusion matrices and croppping feayures predictions, to demonstrate foreground bias. .
  • In Sec. E, we present the dataset information for training and evaluation.
Appendix A Limitations and broader impact
Limitations: Our aim is to develop a region-language alignment model that effectively integrates local visual and semantic features, thereby improving open-vocabulary dense prediction performance. Compared to previous pre-trained Vision-Language Models (VLMs) [4, 55, 71, 59], our proposed DenseVLM achieves superior results and significantly improves downstream task performance. We believe DenseVLM has even greater potential. 1) Scalability. DenseVLM is designed within an efficient, unsupervised region-language alignment framework, making it adaptable to various datasets. However, computational resource limitations have restricted our ability to scale to larger datasets. 2) Model capacity. We employ the ViT-L/14 model from CLIPSelf [59] as a powerful Pre-trained VLM (P-VLM). Utilizing more robust VLMs can yield better performance, and transferring their rich semantic knowledge to training models is a promising direction. 3) Fine-grained Semantic. We categorize objects into broad thing and stuff classes. Fine-grained semantic segmentation and decoupled alignment would enhance the model’s ability to distinguish between similar categories. We plan to explore these avenues in our future research.
@@ Appendix A Limitations and broader impact
warmup steps [14] 1000
@@
learning rate scheduler cosine decay [44]
@@ Appendix B Training details
For open-vocabulary segmentation, we train the models such as SAN [64] and CAT-Seg [5] on the COCO-Stuff [1] dataset for 80kk iterations. For open-vocabulary detection, models are trained for 3 epochs on the OV-COCO [3] benchmark and 48 epochs on OV-LVIS [17] benchmark.
@@ Appendix C Additional experiments
Ablation study on using region proposals. Following RegionCLIP [71] for fine-tuning VLMs with pseudo-labelled region-text pairs, we compare our approach to CLIPSelf [59] in utilizing these pairs. As shown in Tab. 11, CLIPSelf substitutes random image crops with pseudo region-text pairs, resulting in an enhanced recognition for foreground objects while concurrently observing a reduction in the accuracy of background identification. In contrast, our proposed DenseVLM achieves a notable improvement in the recognition accuracy of foreground objects while also improving the identification of background stuff.
@@
Ablation study on various backbones. DenseVLM exhibits adaptability to diverse backbones. As shown in Tab. 13, our models achieve consistent superiority over prior approaches [4, 59] across all dense prediction tasks. Particularly, the ViT-B/16-based DenseVLM performs comparably to the ViT-L/14-based CLIPSelf [59]. Utilizing ViT-L/14 with a large number of parameters as initialization, DenseVLM achieves clearly enhancements across all evaluated metrics, thereby facilitating superior performance in dense prediction tasks.
@@
Table 13: Ablation study on various backbones. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks. * indicates the model initialized by OpenCLIP [4].
@@
Ablation study on different category sets. To assess the impact of varying category sets, we conduct experiments using four various sets: 133 (80), 171 (80), 273 (160), and 316 (160), categorized into foreground and background classes. The set of 133 categories exclusively comprises COCO Panoptic [40] class set, while the 171-category set consists solely of COCO-Stuff [1] class set. The 273-category set integrates non-overlapping classes from both COCO-Stuff and the ADE20K Panoptic [72] dataset, which contains 150 categories. The 316-category set encompasses selected background classes from COCO-Stuff and the ADE20K dataset, which includes 847 categories. Our code accurately reflects all specified category sets.
@@
Appendix D Visualizations
Confusion matrix. We compare the region classification results of our method against previous approaches through a confusion matrix visualization for panoptic masks (both thing and stuff categories) on the COCO Panoptic dataset. These confusion matrixes offer a systematic overview of region classification performance, illustrating the incidence of accurate and erroneous classifications, particularly facilitating a precise assessment of models’ accuracy in differentiating between thing and stuff categories. As shown in Fig. 7, Fig. 8, and Fig. 9, prior methods, including EVA-CLIP [55], RegionCLIP [71], and CLIPSelf [59], often misclassify background regions as co-occurring foreground classes, such as incorrectly identifying snow as skis or sky as kite. In contrast, as demonstrated in Fig. 10, our DenseVLM achieves higher accuracy in recognizing each category, with a notable improvement in the precision of background object identification.
Image grid patches classification. We visualize the classification results of image grid patches using the powerful ViT-L/14 model from CLIPSelf [59]. As shown in Fig. 11, the model focuses heavily on foreground object recognition, but significant portions of background patches are misclassified as foreground objects. The training VLMs are prone to learning these errors. Furthermore, regions with incorrect classifications often have low confidence scores, highlighting the importance of filtering them out.
Appendix E Datasets of training and evaluation
COCO: COCO [40] is a large-scale panoptic segmentation dataset encompassing 80 Thing and 53 Stuff categories. The dataset comprises 118,000 images designated for the training set and 5,000 images for the validation set.
ADE20k: ADE20k [73] spans a broad spectrum of indoor and outdoor scenes, comprising 2,000 images for the validation set. This dataset includes 100 Thing and 50 Stuff categories. We evaluate open-vocabulary semantic annotations using both the extensive 847-category version (referred to as A-847) and the more frequently adopted 150-category version (referred to as A-150).
Pascal Context: Pascal-Context [10] constitutes an extensive dataset derived from Pascal-VOC 2010. We evaluate open-vocabulary semantic annotations using the complete set of 459 classes, referred to as PC-459.
OV-COCO: The open-vocabulary detection COCO (OV-COCO) benchmark, introduced in OV-RCNN [68], divides the 65 object categories in the COCO dataset into 48 base categories and 17 novel categories.
OV-LVIS: The open-vocabulary detection LVIS (OV-LVIS) benchmark, introduced in ViLD [15], defines the 337 rare categories from LVIS v1.0 dataset [17] as novel categories.

### v4

@@ 4.2 Comparison with other VLMs
Quantitative evaluation. We perform a comprehensive quantitative evaluation of dense representations across multiple VLMs on the COCO Panoptic [lin2014microsoft] and ADE20K Panoptic [zhou2017scene] datasets. As shown in Tab. 1, while previous methods [openclip, fang2023data, siglip, sun2023eva] achieve strong zero-shot image classification performance, their performance in region recognition is notably suboptimal. For instance, EVA-CLIP [sun2023eva] achieves only a Top-1 box classification accuracy of 44.3% on COCO and 33.0% on ADE20K. Although RegionCLIP [zhong2022regionclip], which is trained on region-text pairs, shows improved performance on COCO, but fails to generalize to datasets such as ADE20K. Moreover, FineCLIP [jing2025fineclip] and CLIPSelf [wu2023clipself], which incorporate self-distillation, achieve higher Mask-T classification accuracy but perform poorly in Mask-S classification. In contrast, our method, DenseVLM, notably outperforms these models, achieving a 4.3% improvement in Top-1 Mask-T accuracy and a 3.9% improvement in Top-1 Mask-S accuracy on COCO, highlighting its effectiveness in dense prediction tasks.
@@ 4.3 Ablation study
In our ablation study, DenseVLM is trained on the unlabeled images from the COCO train2017 split using $\timesA4$0 GPUs and evaluated on the val2017 split. The experiments in Sec. 4.3 utilize the ViT-B/16 model from EVA-CLIP [sun2023eva] for its superior efficiency and capacity.
Framework design exploration. As shown in Tab. 2, we conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [wu2023clipself] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction while introducing foreground bias issues. In contrast, DenseVLM leverages a more efficient feature-cropping strategy and a decoupled alignment framework to achieve superior performance while substantially reducing both training time and GPU memory usage.
@@
Figure 6: Zero-shot comparisons of models pre-trained on datasets with three different scales. We select three training sets from the SA-1B dataset [kirillov2023segment]: 100K, 1.1M, and 5.5M seen samples and perform the zero-shot evaluation on the COCO and ADE20K benchmarks.
@@
To address the limitations of category set annotations, we generate category sets using the generative model NXTP [Yue_2024_CVPR] based on images and use DeepSeek-R1 [guo2025deepseek] to classify foreground and background categories. As shown in Tab. 4, our model still achieves strong performance.
@@ 4.4 Zero-shot comparisons at different data scales
Table 6: Results and comparisons of different backbones. * indicates the model initialized by OpenCLIP [openclip].
@@
PACL [mukhoti2023open] 31.4 - -
OVSeg [ovseg] 24.8 7.1 11.0
MAFT [jiao2023learning] 29.1 10.1 12.6
SED [xie2023sed] 31.6 11.4 18.6
SCAN [Liu_2024_CVPR] 30.8 10.8 13.2
CAT-Seg+CLIPSelf [wu2023clipself] 29.7 10.1 -
CAT-Seg+FineCLIP [jing2025fineclip] 32.4 12.2 -
SAN [san] 27.4 10.0 13.0
@@ 4.5 Extending to different backbones
To demonstrate the generality of our method, we extend it to several backbones, including ViT-B/16, R50x4, and ConvNeXt-B from OpenCLIP [openclip], as well as ViT-B/16 from EVA-CLIP [sun2023eva]. From Tab. 6, we observe that CLIPSelf [wu2023clipself] is more effective with ViT-based architectures but is less effective with CNN-based architectures, particularly showing degraded performance with R50x4 compared to the baseline. In contrast, our approach yields consistent performance gains across all architectures, thereby broadening its applicability and potential. Furthermore, DenseVLM narrows the performance gap between powerful models like ViT-L/14 fine-tuned by CLIPSelf and lightweight models, with some lightweight models even outperforming the more powerful ones on the Masks-S metric.
@@ 4.6 Application to open-vocabulary dense tasks
VLDet [lin2022learning] RN50 32.0 50.6 45.8
F-VLM [kuo2022f] RN50 28.0 - 39.6
BARON-Cap [Wu_2023_CVPR] RN50 33.1 54.8 49.1
CORA [CORA] RN50 35.1 35.5 35.4
RO-ViT [Kim_2023_CVPR] ViT-B/16 30.2 - 41.5
RO-ViT [Kim_2023_CVPR] ViT-L/16 33.0 - 47.7
F-ViT+CLIPSelf [wu2023clipself] ViT-B/16 25.4 40.9 36.8
F-ViT+FineCLIP [jing2025fineclip] ViT-B/16 29.8 45.9 41.7
F-ViT [wu2023clipself] ViT-B/16 17.5 41.0 34.9
@@
VLDet [lin2022learning] RN50 21.7 29.8 30.1
BARON-Cap [Wu_2023_CVPR] RN50 22.6 27.6 27.6
F-VLM [kuo2022f] RN50 18.6 - 24.2
COR [CORA] RN50x4 22.2 - -
RO-ViT [Kim_2023_CVPR] ViT-B/16 28.0 - 30.2
F-ViT+CLIPSelf [wu2023clipself] ViT-B/16 10.6 7.6 9.3
F-ViT+FineCLIP [jing2025fineclip] ViT-B/16 10.4 8.0 9.5
F-ViT [wu2023clipself] ViT-B/16 11.5 12.3 15.4
@@
Open-vocabulary semantic segmentation. We apply DenseVLM models initialized with OpenAI CLIP [CLIP] to open-vocabulary semantic segmentation, including SAN [san] with a frozen backbone and Cat-Seg [catseg] with a fine-tuned backbone. The models are trained on COCO-Stuff [caesar2018coco] and evaluated on the ADE20K [ade20k] (ADE-150 and ADE-847) and PASCAL Context [pascal_context] (PC-459) datasets, using the mean Intersection-over-Union (mIoU) metric. As shown in Tab. 7, DenseVLM consistently improves performance across all evaluation datasets, further enhancing the state-of-the-art performance.
Open-vocabulary object detection. Building on previous work [wu2023clipself], we construct open-vocabulary object detectors using the F-ViT architecture, a two-stage detector built on frozen ViTs from EVA-CLIP [sun2023eva]. As shown in Tab. 8, we evaluate performance on the OV-COCO [chen2015microsoft] benchmark by reporting box AP at IoU 0.5 for base, novel, and overall categories ($_{50}^{\mathrm{novel}}, AP$$_{50}^{\mathrm{base}}$, AP50), and on OV-LVIS [Gupta_2019_CVPR] benchmark by mean AP for rare (mAPr), common (mAPc), and all categories (mAP). Replacing the frozen EVA-CLIP ViTs with DenseVLM models leads to clear performance improvements on both benchmarks, achieving competitive performance relative to previous methods.
@@ Acknowledgments
  • In Sec. 6, we detail the limitations and broader impact of DenseVLM.
  • In Sec. 7, we present the training details.
  • In Sec. 8, we offer supplementary experiments on input image sizes, region proposals, backbones, category sets and threshold of region denoising for training VLMs.
  • In Sec. 9, we show visualizations, including confusion matrices and croppping feayures predictions, to demonstrate foreground bias. .
  • In Sec. 10, we present the dataset information for training and evaluation.
6 Limitations and broader impact
Limitations: Our aim is to develop a region-language alignment model that effectively integrates local visual and semantic features, thereby improving open-vocabulary dense prediction performance. Compared to previous pre-trained Vision-Language Models (VLMs) [openclip, sun2023eva, zhong2022regionclip, wu2023clipself], our proposed DenseVLM achieves superior results and significantly improves downstream task performance. We believe DenseVLM has even greater potential. 1) Scalability. DenseVLM is designed within an efficient, unsupervised region-language alignment framework, making it adaptable to various datasets. However, computational resource limitations have restricted our ability to scale to larger datasets. 2) Model capacity. We employ the ViT-L/14 model from CLIPSelf [wu2023clipself] as a powerful Pre-trained VLM (P-VLM). Utilizing more robust VLMs can yield better performance, and transferring their rich semantic knowledge to training models is a promising direction. 3) Fine-grained Semantic. We categorize objects into broad thing and stuff classes. Fine-grained semantic segmentation and decoupled alignment would enhance the model’s ability to distinguish between similar categories. We plan to explore these avenues in our future research.
@@ 6 Limitations and broader impact
warmup steps [goyal2018accurate] 1000
@@
learning rate scheduler cosine decay [loshchilov2017sgdr]
@@ 7 Training details
For open-vocabulary segmentation, we train the models such as SAN [san] and CAT-Seg [catseg] on the COCO-Stuff [caesar2018coco] dataset for 80kk iterations. For open-vocabulary detection, models are trained for 3 epochs on the OV-COCO [chen2015microsoft] benchmark and 48 epochs on OV-LVIS [Gupta_2019_CVPR] benchmark.
@@ 8 Additional experiments
Ablation study on using region proposals. Following RegionCLIP [zhong2022regionclip] for fine-tuning VLMs with pseudo-labelled region-text pairs, we compare our approach to CLIPSelf [wu2023clipself] in utilizing these pairs. As shown in Tab. 11, CLIPSelf substitutes random image crops with pseudo region-text pairs, resulting in an enhanced recognition for foreground objects while concurrently observing a reduction in the accuracy of background identification. In contrast, our proposed DenseVLM achieves a notable improvement in the recognition accuracy of foreground objects while also improving the identification of background stuff.
@@
Ablation study on various backbones. DenseVLM exhibits adaptability to diverse backbones. As shown in Tab. 13, our models achieve consistent superiority over prior approaches [openclip, wu2023clipself] across all dense prediction tasks. Particularly, the ViT-B/16-based DenseVLM performs comparably to the ViT-L/14-based CLIPSelf [wu2023clipself]. Utilizing ViT-L/14 with a large number of parameters as initialization, DenseVLM achieves clearly enhancements across all evaluated metrics, thereby facilitating superior performance in dense prediction tasks.
@@
Table 13: Ablation study on various backbones. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks. * indicates the model initialized by OpenCLIP [openclip].
@@
Ablation study on different category sets. To assess the impact of varying category sets, we conduct experiments using four various sets: 133 (80), 171 (80), 273 (160), and 316 (160), categorized into foreground and background classes. The set of 133 categories exclusively comprises COCO Panoptic [lin2014microsoft] class set, while the 171-category set consists solely of COCO-Stuff [caesar2018coco] class set. The 273-category set integrates non-overlapping classes from both COCO-Stuff and the ADE20K Panoptic [zhou2017scene] dataset, which contains 150 categories. The 316-category set encompasses selected background classes from COCO-Stuff and the ADE20K dataset, which includes 847 categories. Our code accurately reflects all specified category sets.
@@
9 Visualizations
Confusion matrix. We compare the region classification results of our method against previous approaches through a confusion matrix visualization for panoptic masks (both thing and stuff categories) on the COCO Panoptic dataset. These confusion matrixes offer a systematic overview of region classification performance, illustrating the incidence of accurate and erroneous classifications, particularly facilitating a precise assessment of models’ accuracy in differentiating between thing and stuff categories. As shown in Fig. 7, Fig. 8, and Fig. 9, prior methods, including EVA-CLIP [sun2023eva], RegionCLIP [zhong2022regionclip], and CLIPSelf [wu2023clipself], often misclassify background regions as co-occurring foreground classes, such as incorrectly identifying snow as skis or sky as kite. In contrast, as demonstrated in Fig. 10, our DenseVLM achieves higher accuracy in recognizing each category, with a notable improvement in the precision of background object identification.
Image grid patches classification. We visualize the classification results of image grid patches using the powerful ViT-L/14 model from CLIPSelf [wu2023clipself]. As shown in Fig. 11, the model focuses heavily on foreground object recognition, but significant portions of background patches are misclassified as foreground objects. The training VLMs are prone to learning these errors. Furthermore, regions with incorrect classifications often have low confidence scores, highlighting the importance of filtering them out.
10 Datasets of training and evaluation
COCO: COCO [lin2014microsoft] is a large-scale panoptic segmentation dataset encompassing 80 Thing and 53 Stuff categories. The dataset comprises 118,000 images designated for the training set and 5,000 images for the validation set.
ADE20k: ADE20k [ade20k] spans a broad spectrum of indoor and outdoor scenes, comprising 2,000 images for the validation set. This dataset includes 100 Thing and 50 Stuff categories. We evaluate open-vocabulary semantic annotations using both the extensive 847-category version (referred to as A-847) and the more frequently adopted 150-category version (referred to as A-150).
Pascal Context: Pascal-Context [everingham2010pascal] constitutes an extensive dataset derived from Pascal-VOC 2010. We evaluate open-vocabulary semantic annotations using the complete set of 459 classes, referred to as PC-459.
OV-COCO: The open-vocabulary detection COCO (OV-COCO) benchmark, introduced in OV-RCNN [Zareian_2021_CVPR], divides the 65 object categories in the COCO dataset into 48 base categories and 17 novel categories.
OV-LVIS: The open-vocabulary detection LVIS (OV-LVIS) benchmark, introduced in ViLD [gu2021open], defines the 337 rare categories from LVIS v1.0 dataset [Gupta_2019_CVPR] as novel categories.